OpenStudy (anonymous):

At the end of the first day, 7 weeds appear in your neighborhood park. Each day, the number of weeds increases by 4 times. How many weeds will be in the park at the end of 14 days? a. 35,456,231 b. 626,349,395 c. 71 d. 392 I thinks b, am i right?

OpenStudy (anonymous):

First day: 7 weeds Second day: 7*4 weeds Third day: 7*4*4 = 7*(4^2) weeds ... => 14th day: 7*(4^13)= 7*67108864 = 469762048